Q8- Find the slope and the intercept point and Double check the answer 5.52 7.67 8.06 9.57 10.55 12.44 1.5 2.5 3.5 Q9- There are two Arrays A and B for those two Arrays find G6 2 1 1 1 2 3 3 -1 3 2 1- The Determent of Array (b) 2- 5*(A+B) 3- A.B-l 4 2 (x+Y) 5- E(Xi Yi)

hi my dear friends.
I need the solution for these equations.
thanks
Answer #1

Q8)

To find the slope in excel, first enter the values in X column and Y column.

Then select all the values and click on "Charts"

Then select "Scatter chart with lines and markers"

A graph will pop up as shown in figure below

1 4 0 86420

Now click on the line in the graph. It must select all the points.

Right click and select "Add trend line"

14 12 10 Series Y Fll Outline Reset to Match Style IChange Series Chart Type... select Data 3-D Rotation... Add Data Labels

In the pop-up check "Display equation on chart"

Forecast Forward Backward Set Intercept Display Equation on chart 0.0 periods 0.0 periods 0.0 Display B-squared value on char

The value before x in the equation is the slope. And the value after + is intercept

B7 12.44 5.52 7.67 8.06 9.57 10.55 12.44 y 2.5571x+3.2148 14 12 10 4 2 3.5 12 1.5 3.5 17

The slope is 2.5571 and the intercept is 3.2148

Q9)

To create the array in excel, we need to first select the number of rows and columns required. Here 3 rows and 3 columns. Select A1, A2, A3, B1, B2, B3, C1, C2, and C3 for array A

Then we need to enter the formula ={1,3,2;2,3,-1;-2,1,2} and CTRL+SHIFT+ENTER to create the array A

Similarly for array B select E1 to G3 and enter the formula ={1,-1,3;2,-1,1;-1,2,3}

xf1,-1,3;2,-1,1;-1,2,3) 1 2 -2 1 2 1 2

We can also name the array by clicking on "formula" tab and then click on "Define name" enter the name of array and click on ok

1- To find the dterminent of B: we need to use the formula =MDETERM(E1:G3)

H1 X-MDETERM(E1:G3)H 2 2 2 -2 2 2

2- First find (A+B) using the formula =SUM(A1:C3*E1:G3) and CTRL+SHIFT+ENTER

Then the formula =5*I1 and press enter

X-SUM(A1:C3 E1:G3)H 1 1A 1 70 -2 1

3- First we need to transpose the array B. This is done by using the formula =TRANSPOSE(E1:G3) and CRTL+SHIFT+ENTER

J1 丘 11 {=TRANSPOSE(E1:G3)) 《i 21 2 1 14 2 2 7011 1 2 2The to multiply two arrays, we can use =SUMPRODUCT(A1:C3,J1:L3) and CRTL+SHIFT+ENTER

M1 x ﹀ 女 {:SUMPRODUCT(A1:C3,J113)) 1 -1311 14 1 2-11·1 1 1

4- First find the square of arrays using the formula ={1,3,2;2,3,-1;-2,1,2}^2? for array A and enter CTRL+SHIFT+ENTER

(-(1,3,2:2,3,-1;-2,1,212 1 1 1 1 fC) -1 1 4?Similarly Square of array B

E5 {s(1,-1,3;2,-1,1,-1,2,3)^2) -2 4 1 4

Now we can find sum of these two arrays using =SUM(N1:P3*E5:G7)

The answer is 116

5- The formula is =SUM(1*1,3*-1,2*3,2*2,3*-1,-1*1,-2*-1,1*2,2*3)

The answer is 14